OpenStudy (meehan98):

Identify an equation in standard form for an ellipse with its center at the origin, a vertex at (0, 12), and a focus at (0, 8).

OpenStudy (meehan98):

It appears that it's a horizontal ellipse so I know that the equation will be \[\frac{ x ^{2} }{ a ^{2} } + \frac{ y ^{2} }{ b ^{2} }=1\]

OpenStudy (anonymous):

That is right.

OpenStudy (meehan98):

I just don't know how to find the other vertex in order to find b^2.

OpenStudy (anonymous):

Actually isn't it a verticle ellipse.

OpenStudy (anonymous):

Because the vertex (furtherest point) is 12 up from the origin

OpenStudy (meehan98):

Oh yes, that's right!

OpenStudy (anonymous):

So the semi major axis is 12, in the y direction

OpenStudy (anonymous):

Therefore the equation is (y^2)/(12^2) + (x^2)/(b^2) = 1

OpenStudy (anonymous):

Now for the b value I think you use the rule that for any point on the ellipse, the sum of the distance from both foci is 2a (a is the semi major axis)

OpenStudy (anonymous):

Ill draw a graph to show

OpenStudy (meehan98):

So, you use the constant sum of an ellipse formula: d=PF1 + PF2

OpenStudy (anonymous):

 in the picture i took the point to be the semi minor axis. Since the ellipse is symmertrical (both foci same distance) the distance from that point to each foci is a (semi major axis)